# Bifurcation mechanism of quasihalo orbit from Lissajous Orbit in the restricted three-body problem

### v1

@@ 4.1 Third-order analytical solution
The solution branch corresponds to northern halo orbits (Class I) for $\eta>0$ and southern halo orbits (Class II) for $\eta<0$. These results align with the classical outcomes of the third-order analytical solution of halo orbits presented in richardson1980analytic . To sum up, quasihalo orbits bifurcate from Lissajous orbits when $\eta$ is a solution of (29). In particular, halo orbits bifurcate from planar Lyapunov periodic orbits when $\eta$ is a solution of (29) with $\beta$ = 0.
Remark 4. The third-order solution (26) serves as an initial approximation to the center manifolds of around the collinear libration points in the RTBP. A higher-order series solution is required for real space missions. In such case, the bifurcation equation is no longer a hyperbola. Figure 1 just provides a basic outline of the feasible region for $\eta$, and in fact, its real feasible region is more intricate than depicted in Fig. 1. A higher-order numerical feasible region of $\eta$ will be presented in the following section.
Figure 2: Feasible region of coupling coefficient $\eta$ for center manifolds of $L_{1}$ in the Earth-Sun system with different orders $n$ = 3, 5, 7, 11, 15, 19. (In this case, $\mu$ = 3.040423398444176e-6, $\gamma_{1}$ = 1.00109772277814e-2) with $\alpha$ and $\beta$.
Figure 3: Contour maps of coupling coefficient $\eta$ within feasible region for center manifolds of $L_{1}$ in the Earth-Sun.
@@ 4.2 Numerical results
The semi-analytical computation of center manifolds up to a certain order $n$ for the given system parameter $\mu$ is implemented to verify the accuracy of the proposed method. B shows the coefficients of expansion for center manifolds of $L_{1}$ in Sun-Earth system, up to order 3.
@@
The third-order analytical solution in Subsection 4.1 illustrates that a feasible region of coupling coefficient $\eta$ of center manifolds is bounded by a hyperbola. Similar to the process in (30) for the third-order solution, higher-order $\Delta(\alpha,\beta)$ can be obtained during the computation of higher-order semi-analytical solutions. For each given paired amplitudes $(\alpha,\beta)$, non-trivial values of $\eta$ can be determined from $\Delta(\alpha,\beta)=0$. Then, The feasible region of $\eta$ of center manifolds around $L_{1}$ in the Earth-Sun system with different orders is presented in Fig. 2, where only the range of $\alpha\in[0,0.35]$, $\beta\in[0,0.4]$ and $\eta\in(0,3.0]$ is considered due to the symmetry of $\eta$ and the divergence of (17) for large amplitudes. The colorbar represents the number of solutions $N(\eta)$ of $\Delta(\alpha,\beta)=0$ with the range $\eta\in(0,3.0]$. It can be seen that $N$ has only two possible values $N=0$ or $N=1$, distributed on both sides of a hyperbola boundary when the order $n=3$. This result is consistent with the analytical solution in Subsection 4.1. As the order increases, parts of the region with $N=0$ and $N=1$ are replaced by the region with $N(\eta)>1$. This means that, with increasing amplitudes $\alpha$ and $\beta$, more than one quasihalo orbit bifurcates from a Lissajous orbit.
Figure 4: Periodic and quasi-periodic orbits around $L_{1}$ in the Earth-Sun system and its projection on $xy$, $xz$, and $yz$ planes, red: planar Lyapunov orbit with $\alpha$ = 0.05, $\beta$ = 0.0; black: vertical Lyapunov orbit with $\alpha$ = 0.0, $\beta$ = 0.25; blue: Lissajous orbit with $\alpha$ = 0.05, $\beta$ = 0.25.
Figure 5: Quasi-periodic orbits around $L_{1}$ in the Earth-Sun system with $\alpha$ = 0.144227 and $\beta$ = 0.04 and its projection on $xy$, $xz$, and $yz$ planes. (a) $\eta$ = 0, Lissajous orbit (b) $\eta$ = 0.08180669069, quasihalo orbit.
@@
Remark 5. The polynomials $\Delta(\alpha,\beta)$ may encompass all the information about local bifurcation around collinear libration points in the RTBP. It is reasonable to believe that with further increases in order, the number of solutions of $\eta$ will also increase. Furthermore, it is observed that the number of solutions for $\eta$ multiplies (due to only $\eta\in(0,3.0]$ being considered, some solutions for $\eta$ are omitted here). This is a typical period-doubling bifurcation, indicating that chaos naturally occurs in the center manifolds of collinear libration points in the RTBP. However, the high-order series solution (17) cannot describe this phenomenon due to its divergence.
@@
Then, a contour map of $\eta$ values with an order $n=19$ is computed as shown in Fig. 3. The feasible region corresponds to the bottom-right part of Fig. 2. Each contour map describes the distribution of one solution of $\eta$. Empty regions indicate no solution for $\eta$. Six contour maps represent a maximum of six solutions of $\eta$ at the 19-th order. Figure 3 shows that the feasible region is largest when there is only one non-zero solution for $\eta$ (top left). Subsequently, as the number of non-zero $\eta$ solution increases, the feasible region becomes small. When the amplitude $\alpha$ and $\beta$ are small, there are no non-zero $\eta$, corresponding to the empty region in the top-left part of Fig. 3. In this case, the higher-order series (17) only describes planar and vertical Lyapunov orbits and Lissajous orbits. Figure 4 shows a typical plot of planar Lyapunov orbits, vertical Lyapunov orbits and Lissajous orbits in the synodic coordinate system.
@@
Upon increasing $\alpha$ and $\beta$ to the region with one solution of $\eta$, Fig. 3 shows that $\eta$ values change from zero to non-zeros but they are very small, which means the week coupling of two degrees of freedom (in-plane and out-of-plane motions). In this case, quasihalo orbits closely resemble the corresponding Lissajous orbits, with the amplitudes identical to quasihalo orbits but $\eta$ values being zeros. This observation is evident from Fig. 5, where $\eta$ = 0.08180669069. As the coupling effect ($\eta$ 0) of the in-plane motion acts on the out-of-plane motion, a plane-symmetric Lissajous orbit (linear part in $z$-direction: $z_{1}=\beta\cos\theta_{2}$) becomes an approximately oblique-upward symmetric quasihalo orbit (linear part in $z$-direction: $z_{1}=\beta\cos\theta_{2}+\eta\alpha\cos\theta_{1}$).
@@
With further increment in $\alpha$ and $\beta$, one $\eta$ value rapidly grows larger and more solutions for $\eta$ appear. Figure 7 shows two quasihalo orbits with the same amplitudes. It can be seen that the quasihalo orbit with the smaller value of $\eta$ exhibits week coupling and remains similar to Lissajous orbits. The larger value of $\eta$ leads to a significant coupling effect on the motion in z-directions, resulting in a typical quasihalo orbit, as shown in Fig. 7b. In this case, in-plane and out-of-plane motions are strong coupling. As $\alpha$ and $\beta$ increase to the region shown in the bottom-left of Fig.3, four solutions for $\eta$ can be found from $\Delta=0$, i.e., $N(\eta)=4$. Figure 8 shows four quasihalo orbits corresponding to these four solutions for the given $\alpha$ and $\beta$. It can be seen they have some strange bendings, likely due to the corresponding $\alpha$ and $\beta$ not being within the practical region of convergence.
@@
Figure 6 shows the domain of practical convergence of the proposed analytical solution up to order 35 for Lissajous orbits ($\eta=0$) and Quasihalo orbits ($\eta\neq 0$) around $L_{1}$ of the Earth-Sun system. Here Fig. 6(a) is similar the result in jorba1999dynamics . This is because, when $\eta$ is zero, no bifurcation occurs, and the analytical solution is identical to the solution of Lissajous orbits presented in jorba1999dynamics . However, when $\eta$ is non-zero, coupling effects between different degrees of freedom lead to the bifurcation, specifically the generation of quasihalo orbits. Figure 6(b), for the first time, provides the actual convergence domain of the approximate analytical solution for quasihalo orbits generated by the first bifurcation from Lissajous orbits. It can be observed that the practical convergence domain for quasihalo orbits is significantly smaller than the practical convergence domain for the Lissajous orbit. This is due to the coupling in the plane direction causing a large actual amplitude in the z-direction, even when the amplitude $\beta$ is small.
Figure 6: Domain of practical convergence of the proposed analytical solution up to order $n$ = 35 for (a) Lissajous orbits and (b) Quasihalo orbits around $L_{1}$ of the Earth-Sun system
@@
It is known that center manifolds around collinear libration points in the RTBP are four-dimensional, making direct graphical display challenging. As a byproduct of the proposed analytical method for constructing center manifolds, the dynamics inside center manifolds can be globally described in a two-dimensional Poincaré section by fixing $z=0$ with $z>0$ and the Jacobian integral $C=C_{0}$ in synodic coordinate system. First, arbitrary phase angles $\phi_{1}$ and $\phi_{2}$ and arbitrary initial time t0 from (17) is chosen. For the chosen $C_{0}$, paired amplitudes ($\alpha$, $\beta$) and corresponding initial states ($x_{0}y_{0}z_{0}$) can be computed. Then, starting from each initial state, every points $(x_{i},y_{i})$ is plotted in the two-dimensional Poincaré section when $z_{i}=0$ within a specified time interval according to (17).
Figure 7: Two quasihalo orbits around $L_{1}$ in the Earth-Sun system with $\alpha$ = 0.167 and $\beta$ = 0.055 and its projection on $xy$, $xz$, and $yz$ planes. (a) $\eta$ = 0.04676813553 (b) $\eta$ = 1.552696086.
Figure 8: Four quasihalo orbits around $L_{1}$ in the Earth-Sun system with $\alpha$ = 0.180 and $\beta$ = 0.142 and its projection on $xy$, $xz$, and $yz$ planes. (a) $\eta$ = 0.9190215098 (b) $\eta$ = 1.12844267 (c) $\eta$ = 1.636802107 (d) $\eta$ = 1.870762835.
@@
Figure 9 shows two Poincaré sections with two different Jacobian integrals, defining a closed region. The boundary of the region is a planar Lyapunov orbit with the paired amplitudes ($\alpha_{\rm{max}}$, 0) and a fixed point on $y$-axis represents a vertical Lyapunov orbit with the paired amplitudes (0, $\beta_{\rm{max}}$), where $\alpha_{\rm{max}}$ and $\beta_{\rm{max}}$ are maximum in-plane and out-of-plane amplitudes for the selected $C_{0}$, respectively. Other circles inside this region respond to Lissajous orbits with the paired amplitudes ($\alpha$, $\beta$) where $\alpha$ $\alpha_{\rm{max}}$ and $\beta<\beta_{\rm{max}}$. The bifurcation occurs when the increase in $C$ results in a non-zero solution $\eta\neq 0$ for the equation $\Delta=\sum\limits_{0\leq i+j\leq n}{{d_{ij}}{\alpha^{i}}{\beta^{j}}=0}$. This visual computing not only shows the emergence of well-known halo orbits from planar Lyapunov orbits but also illustrates the generation of quasi-periodic orbits from Lissajous orbits.
Figure 9: Poincaré section of the center manifold of $L_{1}$ in the Earth-Sun system: (a) $C$ = 3.000842. (b) $C$ = 3.000804.
Remark 6. The computed results presented in Fig. 9 are similar to those found in previous studies (gomez2001dynamics, ; jorba1999dynamics, ). However, these results were computed through numerical continue methods with high computational complexity or displayed in normal coordinate via a complicated normal form. In contrast, the Poincaré sections of center manifolds can be directly derived from the high-order series solution (17) in this paper.
@@ 5 Conclusions
Understanding the coupling between in-plane and out-of-plane motions and its connection to the bifurcation of halo/quasihalo orbits is crucial for unraveling the mechanism governing their generation and obtaining their unified an analytical solution. This paper introduced a novel concept of coupling coefficients into the RTBP for the first time, incorporated a new correction term into the RTBP equation to characterize this coupling effect and successfully derive a unified analytical solution for the center manifolds of collinear libration points in the RTBP. When the bifurcation equation $\Delta=0$ has no real solutions, the series solution (17) describes planar Lyapunov orbits, vertical Lyapunov orbits and Lissajous orbits. When the equation $\Delta=0$ has only zero solutions, the bifurcation occurs. When the equation $\Delta=0$ yields non-zero real solution, the series solution (17) further describes both halo orbits and quasihalo orbits, with the latter being constructed analytically for the first time. Remarkably, larger amplitudes lead to multiple solutions of the equation $\Delta=0$ with the higher-order series solution, indicating the discovery of a multiple bifurcation, which can be explicitly calculated.
Although the proposed analytical method is aimed at constructing center manifolds around collinear libration points in the RTBP, it can be extended to invariant manifolds in the RTBP and even general multi-degree-of-freedom dynamical systems. Besides, the coupling efficient $\eta$ and the equation $\Delta=0$ serve as conditions for bifurcation. Thus, this method can also be referred to as an analytical bifurcation calculation method.
Appendix A Coefficients of third-order analytical solution
@@
Appendix B Coefficient of expansion for center manifolds of $L_{1}$ in Sun-Earth system (up to order 3)

### v2

@@ 4.1 Third-order analytical solution
The solution branch corresponds to northern halo orbits (Class I) for $\eta>0$ and southern halo orbits (Class II) for $\eta<0$. These results align with the classical outcomes of the third-order analytical solution of halo orbits presented in [14]. To sum up, quasihalo orbits bifurcate from Lissajous orbits when $\eta$ is a solution of (30). In particular, halo orbits bifurcate from planar Lyapunov periodic orbits when $\eta$ is a solution of (30) with $\beta$ = 0.
The third-order solution (26) serves as an initial approximation to the center manifolds of around the collinear libration points in the RTBP. A higher-order series solution is required for real space missions. In such case, the bifurcation equation is no longer a hyperbola. Figure 1 just provides a basic outline of the feasible region for $\eta$, and in fact, its real feasible region is more intricate than depicted in Fig. 1. A higher-order numerical feasible region of $\eta$ will be presented in the following section.
Figure 2: Feasible region of $\eta$ for center manifolds of $L_{1}$ in the Sun-Earth system.
Figure 3: Distribution of $\eta$ within feasible region for center manifolds of $L_{1}$.
@@ 4.2 Numerical results
The semi-analytical computation of center manifolds up to a certain order $n$ for the given system parameter $\mu$ is implemented to verify the proposed method. The analysis includes a thorough examination of the high-order bifurcation process that gives rise to quasihalo orbits from Lissajous orbits. Furthermore, the practical convergence domain of the proposed analytical solution for both Lissajous orbits and quasihalo orbits is provided. In addition, for validation purposes, the coefficients of the expansion up to the third order for center manifolds of $L_{1}$ in Sun-Earth system ($\mu$ = 3.040423398444176×$\times 10^{-6}$, $\gamma_{1}$ = 1.00109772277814×$\times 10^{-2}$) are shows in Appendix 5.2.
@@
The third-order analytical solution in Subsection 4.1 illustrates that a feasible region of coupling coefficient $\eta$ of center manifolds is bounded by a hyperbola. Similar to the process in (30) for the third-order solution, higher-order $\Delta(\alpha,\beta)$ can be obtained during the computation of higher-order semi-analytical solutions. For each given paired amplitudes $(\alpha,\beta)$, non-trivial values of $\eta$ can be determined from $\Delta(\alpha,\beta)=0$. Then, the feasible region of $\eta$ of center manifolds around $L_{1}$ in the Sun-Earth system with different orders is presented in Fig. 2, where only the range of $\alpha\in[0,0.35]$, $\beta\in[0,0.4]$ and $\eta\in(0,3.0]$ is considered due to the symmetry of $\eta$ and the divergence of (17) for large amplitudes. The colorbar represents the number of solutions $N(\eta)$ of $\Delta(\alpha,\beta)=0$ with the range $\eta\in(0,3.0]$. It can be seen that $N$ has only two possible values $N=0$ or $N=1$, distributed on both sides of a hyperbola boundary when the order $n=3$. This result is consistent with the analytical solution in Subsection 4.1. As the order increases, parts of the region with $N=0$ and $N=1$ are replaced by the region with $N(\eta)>1$. This means that, with increasing amplitudes $\alpha$ and $\beta$, the Lissajous orbit undergoes multiple bifurcations, forming multiple families of quasihalo orbits.
Figure 4: Typical orbits around $L_{1}$ in the Sun-Earth system.
@@
The polynomials $\Delta(\alpha,\beta)$ may encompass all the information about local bifurcation around collinear libration points in the RTBP. It is reasonable to believe that with further increases in order, the number of solutions of $\eta$ will also increase. Furthermore, it is observed that the number of solutions for $\eta$ multiplies (due to only $\eta\in(0,3.0]$ being considered, some solutions for $\eta$ are omitted here). However, larger values of $\eta$ might lead to divergence of the solution. This is because the power of $\eta$ in the coefficients of the $n$-th order terms in the series solution (17) does not exceed $2n$. To ensure the convergence of the series solution, it is required that $\eta^{2n}(\alpha^{i}\beta^{j})<1\,(i+j<n)$, which implies that $\eta$ needs to satisfy $\left|\eta\right|<\sqrt{1/{{\{\alpha,\beta\}}_{\max}}}$ at minimum, where $\{\alpha,\beta\}_{\max}$ denotes taking the maximum value between $\alpha$ and $\beta$.
@@
Then, distributions of coupling coefficient $\eta$ values with an order $n=19$ is computed as shown in Fig. 3. The feasible region corresponds to the bottom-right part of Fig. 2. Each contour map describes the distribution of one solution of $\eta$. Empty regions indicate no solution for $\eta$. Six contour maps represent a maximum of six solutions of $\eta$ at the 19-th order. Figure 3 shows that the feasible region is largest when there is only one non-zero solution for $\eta$ (top left). Subsequently, as the number of non-zero $\eta$ solution increases, the feasible region becomes small. When the amplitude $\alpha$ and $\beta$ are small, there are no non-zero $\eta$, corresponding to the empty region in the top-left part of Fig. 3. In this case, the higher-order series (17) only describes planar and vertical Lyapunov orbits and Lissajous orbits. Figure 4 shows a typical plot of planar Lyapunov orbits by the red line ($\alpha$ = 0.05, $\beta$ = 0.0), vertical Lyapunov orbits by the black line ($\alpha$ = 0.0, $\beta$ = 0.25) and Lissajous orbits by the blue line ($\alpha$ = 0.05, $\beta$ = 0.25) in the synodic coordinate system.
@@
Upon increasing $\alpha$ and $\beta$ to the region with one solution of $\eta$, Fig. 3 shows that $\eta$ values change from zero to non-zero but they are very small, which means the weak coupling of two degrees of freedom (in-plane and out-of-plane motions). In this case, quasihalo orbits closely resemble the corresponding Lissajous orbits, with the amplitudes identical to quasihalo orbits but $\eta$ values being zeros. This observation is evident from Fig. LABEL:fig:Lissajous_and_quasihalo_orbit, where $\eta$ = 0.08180669069. As the coupling effect ($\eta$ > 0) of the in-plane motion acts on the out-of-plane motion, a plane-symmetric Lissajous orbit (linear part in $z$-direction: $z_{1}=\beta\cos\theta_{2}$) becomes an approximately oblique-upward symmetric quasihalo orbit (linear part in $z$-direction: $z_{1}=\beta\cos\theta_{2}+\eta\alpha\cos\theta_{1}$).
@@
With further increment in $\alpha$ and $\beta$, one $\eta$ value rapidly grows larger and more solutions for $\eta$ appear. Figure LABEL:fig:Two_quasihalo_orbits shows two quasihalo orbits with the same amplitudes. It can be seen that the quasihalo orbit with the smaller value of $\eta$ exhibits weak coupling and remains similar to Lissajous orbits. The larger value of $\eta$ leads to a significant coupling effect on the motion in the $z$-direction, resulting in a typical quasihalo orbit, as shown in Fig. LABEL:fig:Two_quasihalo_orbitsb. In this case, in-plane and out-of-plane motions are strongly coupled. As $\alpha$ and $\beta$ increase to the region shown in the bottom-left of Fig.3, four solutions for $\eta$ can be found from $\Delta=0$, i.e., $N(\eta)=4$. Figure LABEL:fig:Four_quasihalo_orbits shows four quasihalo orbits corresponding to these four solutions for the given $\alpha$ and $\beta$. It can be seen they have some strange bendings, likely due to the corresponding $\alpha$ and $\beta$ not being within the practical region of convergence.
@@
Figure LABEL:fig:Convergence_domain shows the domain of practical convergence of the proposed analytical solution up to order 35 for Lissajous orbits ($\eta=0$) and quasihalo orbits ($\eta\neq 0$) around $L_{1}$ of the Sun-Earth system, where color scale represent the normalized Euclidean norm of difference in position vectors between the analytical and numerically integrated solutions. Here Fig. LABEL:fig:Convergence_domain(a) is similar to the result in [4]. This is because, when $\eta$ is zero, no bifurcation occurs, and the analytical solution is identical to the solution of Lissajous orbits presented in [4]. However, when $\eta$ is non-zero, coupling effects between different degrees of freedom lead to the bifurcation, specifically the generation of quasihalo orbits. Figure LABEL:fig:Convergence_domain(b), for the first time, provides the actual convergence domain of the approximate analytical solution for quasihalo orbits generated by the first bifurcation from Lissajous orbits. It can be observed that the practical convergence domain for quasihalo orbits is significantly smaller than the practical convergence domain for the Lissajous orbit. This is due to the coupling in the plane direction causing a large actual amplitude in the z-direction, even when the amplitude $\beta$ is small.
@@
Finally, let us visually illustrate the Poincaré section of the center manifold in synodic coordinate system using the solution (17). It is known that center manifolds around collinear libration points in the RTBP are four-dimensional, making direct graphical display challenging. As a byproduct of the proposed analytical method for constructing center manifolds, the dynamics inside center manifolds can be globally described in a two-dimensional Poincaré section by fixing $z=0$ with $z>0$ and the Jacobian integral $C=C_{0}$ in synodic coordinate system. First, arbitrary phase angles $\phi_{1}$ and $\phi_{2}$ and arbitrary initial time $t_{0}$ from (17) is chosen. For the chosen $C_{0}$, paired amplitudes ($\alpha$, $\beta$) and corresponding initial states ($x_{0},y_{0},z_{0}$) can be computed. Then, starting from each initial state, every points $(x_{i},y_{i})$ is plotted in the two-dimensional Poincaré section when $z_{i}=0$ within a specified time interval according to (17).
@@
Figure LABEL:fig:Poincare_section shows two Poincaré sections projected onto the $xy$-plane of the synodic coordinate system. Each section defies a closed region. The boundary of the region is a planar Lyapunov orbit with the paired amplitudes ($\alpha_{\rm{max}}$, 0) and a fixed point on $y$-axis represents a vertical Lyapunov orbit with the paired amplitudes (0, $\beta_{\rm{max}}$), where $\alpha_{\rm{max}}$ and $\beta_{\rm{max}}$ are maximum in-plane and out-of-plane amplitudes for the selected $C_{0}$, respectively. Other circles inside this region respond to Lissajous orbits with the paired amplitudes ($\alpha$, $\beta$) where $\alpha$ < $\alpha_{\rm{max}}$ and $\beta<\beta_{\rm{max}}$. The bifurcation occurs when the increase in $C$ results in a non-zero solution $\eta\neq 0$ for the equation $\Delta=\sum\limits_{0\leq i+j\leq n}{{d_{ij}}{\alpha^{i}}{\beta^{j}}=0}$. This visual computing not only shows the emergence of well-known halo orbits from planar Lyapunov orbits but also illustrates the generation of quasi-periodic orbits from Lissajous orbits.
@@ 5 Conclusions
Understanding the coupling between in-plane and out-of-plane motions and its connection to the bifurcation of halo/quasihalo orbits is crucial for unraveling the mechanism governing their generation and obtaining their unified analytical solution. This paper introduced a novel concept of coupling coefficients into the RTBP for the first time, incorporated a new correction term into the RTBP equation to characterize this coupling effect and successfully derived a unified analytical solution for the center manifolds of collinear libration points in the RTBP. When the bifurcation equation $\Delta=0$ has no real solutions, the proposed series solution describes planar Lyapunov orbits, vertical Lyapunov orbits and Lissajous orbits. When the equation $\Delta=0$ has only zero solutions, the bifurcation occurs. When the equation $\Delta=0$ yields non-zero real solution, the proposed series solution further describes both halo orbits and quasihalo orbits, with the latter being constructed analytically for the first time. Remarkably, larger amplitudes lead to multiple solutions of the equation $\Delta=0$ with the higher-order series solution, indicating the discovery of multiple bifurcations of quasihalo orbits. These solutions can all be explicitly calculated, although some may fall outside the domain of practical convergence due to the order limitation. Finally, it is noted that the coupling efficient $\eta$ and the equation $\Delta=0$ serve as conditions for bifurcation. Thus, this method can also be referred to as an analytical bifurcation calculation method.
Appendix
5.1 Coefficients of third-order analytical solution
@@
5.2 Coefficient of expansion for center manifolds of $L_{1}$ in Sun-Earth system (up to order 3)